A common lipid in brain membranes is called L-alpha-phosphatidylcholine, or PC. PC has a formula of (C subscript 42 H subscript 82 NO subscript 8 P). Each PC molecule occupies an area of 55 square angstroms on one side of the surface of a vesicle. You have 55.0 g of PC, which form into spherical vesicles with a radius of 200. nm. How many vesicles can you form?
1
Expert's answer
2012-10-08T10:00:52-0400
C42H82NO8P have molar mass 759

55.0 n
C42H82NO8P--------------C42H82NO8P
759 & 6.02x10^23

n=4.36 x 10^22 molecules

Square of this molecules is 55 A^2 *4.36 x 10^22 =2.399 x 10^24

Square of vesicles = = 4 pi r^2 = 4*3.14 *200^2 =502400 nm = 5024000 A (1 nm =10 A)

n=2.399 x10^24 / 5024000 =4.775 x 10^17 vesicles